The physician orders an intravenous infusion of 5% dextrose in 0.25% normal saline solution to be infused at 9ml/kg/hour for a 12 month-old infant. The infant weights 29 lbs. How many ml/hr should the nurse infuse of the ordered solution?
Answer #1

Weight of the baby = 29 * 0.454= 13.17 kg

Dose for the baby= 9 * 13.17 = 118.53 ml/hour